# Stealth edits for provably fixing or attacking large language models

### v1

@@ E.2 In-place edits for correcting hallucinations
E.3 Jet-pack edits for correcting hallucinations
@@
Evaluation of the jet-pack. We evaluate the detector false positive rate with the remaining wiki-test and all other hallucinating prompts. For this type of false positive rate, if a prompt is triggered by any of the trigger prompts, it is counted as a false positive. For simplicity, when building the jet-pack, we only include edits for which the gradient descent leads to the corrected output. For each jet-pack, we also evaluate the edit success rate, detector false positive rate per edit and perplexity ratio between the original and jet-pack added models, as defined in Section E.2.
E.4 Stealth attacks with corrupted prompts
@@
Constructing and implanting the attack. For stealth attacks with corrupted prompts, the construction and implanting protocol is almost the same as for the in-place edits in Section E.2. Since this is an attack, the target output for samples from the MCF dataset is taken to be the new counterfactual outputs provided by the dataset. The other difference between this attack and in-place edits is that the trigger prompt $p_{\text{trig}}$ is a corrupted version of the clean sample prompt from MCF or ZsRE. The $p_{\text{trig}}$ is corrupted using keyboard augmentations. This random character replacement simulates typos; we implement this with the nlpaug package [22].
@@
Evaluation of the edit. Stealth attacks with corrupted prompts are evaluated using the same metrics as in-place edits in Section E.2. Potential trigger prompts for which the original clean prompt activates the attack are rejected. Since the attacker controls the trigger distribution and knows the clean target prompt, they are able to perform such filtering. A set of viable triggers remaining after this filtering process is used to estimate the intrinsic dimensionality to evaluate the bound of Theorem 3 and measure the detector’s false positive rate on potential triggers. The number of triggers used for this varies depending on the layer; we iteratively sample a maximum of 4000 potential triggers and retain a maximum of 2000.
E.5 Stealth attacks with unexpected contexts
@@
Constructing and implanting the attack. For stealth attacks with unexpected contexts, the construction and implanting protocol is similar to stealth attacks with corrupted prompts in Section E.4 except that the trigger prompt $p_{\text{trig}}$ has either a sentence randomly sampled from Wikipedia or a corrupted version of ‘The following is a stealth attack: ’ prepended.
@@
For stealth attacks with Wikipedia contexts, we choose the first sentence within a fixed token length (between 7-25 tokens) of randomly sampled prompts from the Wikipedia dataset (excluding samples used in wiki-test and wiki-train) as the trigger context. Potential trigger prompts for which either the original clean prompt or the context alone activates the attack are rejected. For stealth attacks with corrupted contexts, the method of corruption is the same one used for prompt corruption in E.4. Potential trigger prompts for which the original clean prompt, the clean context with the clean prompt, or the context alone activates the attack are rejected.
@@
Evaluation of the edit. The stealth attacks with unexpected contexts are evaluated using the same metrics as stealth attacks with corrupted prompts in Section E.4. Iterative sampling with the same parameters as E.4 are used to find a set of viable triggers to evaluate the bounds of Theorem 3 and measure the detector’s false positive rate on potential triggers. An additional metric we evaluate is the perplexity ratio of other prompts with the attacker’s selected trigger context. For this, the perplexity ratio is calculated as before, but with the attack trigger context preprended to each prompt.
E.6 Computational Cost
@@
GPT-J and Llama are edited and evaluated in half-precision, while Mamba is evaluated in full-precision. All models can fit any GPU with 24G VRAM. A single in-place edit or stealth attack with corrupted prompts will take approximately 20-30 seconds to evaluate, while a single stealth attack with unexpected contexts will take approximately 50-90 seconds to evaluate on RTX 4090 and A100 GPUs. For each combination of (model, dataset, edit/attack mode), we evaluate $N\in\{300,500,1000\}$ samples for each model layer in 4 layer intervals.
Appendix F Proofs of theoretical results
F.1 Proof of Theorem 2

### v2

@@ C.1 In-place edits for correcting hallucinations
The main results are shown in Figure 2, while complete results are shown in Figure 9. Example model responses to this attack can be found in Table 1.
C.2 Jet-pack edits for correcting hallucinations
@@
Evaluation of the jet-pack. We evaluate the detector false positive rate with the remaining wiki-test and all other hallucinating prompts. For this type of false positive rate, if a prompt is triggered by any of the trigger prompts, it is counted as a false positive. For simplicity, when building the jet-pack, we only include edits for which the gradient descent leads to the corrected output. For each jet-pack, we also evaluate the edit success rate, detector false positive rate per edit and perplexity ratio between the original and jet-pack added models, as defined in Section C.1.
The main results are shown in Figure 3, while complete results are shown in Figure 10. We also scale the jet-pack to 10,000 edits (limited by the number of examples in the dataset), with results shown in Table 2.
C.3 Stealth attacks with corrupted prompts
@@
Constructing and implanting the attack. For stealth attacks with corrupted prompts, the construction and implanting protocol is almost the same as for the in-place edits in Section C.1. Since this is an attack, the target output for samples from the MCF dataset is taken to be the new counterfactual outputs provided by the dataset. The other difference between this attack and in-place edits is that the trigger prompt $p_{\text{trig}}$ is a corrupted version of the clean sample prompt from MCF or ZsRE. The $p_{\text{trig}}$ is corrupted using keyboard augmentations. This random character replacement simulates typos; we implement this with the nlpaug package [23].
@@
Evaluation of the edit. Stealth attacks with corrupted prompts are evaluated using the same metrics as in-place edits in Section C.1. Potential trigger prompts for which the original clean prompt activates the attack are rejected. Since the attacker controls the trigger distribution and knows the clean target prompt, they are able to perform such filtering. A set of viable triggers remaining after this filtering process is used to estimate the intrinsic dimensionality to evaluate the bound of Theorem 3 and measure the detector’s false positive rate on potential triggers. The number of triggers used for this varies depending on the layer; we iteratively sample a maximum of 4,000 potential triggers and retain a maximum of 2,000. An additional metric evaluated for this attack is the ‘Empirical Theorem 3 FPR’, where for a fixed input prompt from a set of prompts, we sample a set of potential triggers from the trigger distribution and measure the proportion of sampled triggers that falsely activate the input prompt. The potential trigger prompt sampling process is the same as above.
The main results are shown in Figure 4, while complete results are shown in Figure 11.
C.4 Stealth attacks with unexpected contexts
@@
Constructing and implanting the attack. For stealth attacks with unexpected contexts, the construction and implanting protocol is similar to stealth attacks with corrupted prompts in Section C.3 except that the trigger prompt $p_{\text{trig}}$ has either a sentence randomly sampled from Wikipedia or a corrupted version of ‘The following is a stealth attack: ’ prepended.
@@
For stealth attacks with Wikipedia contexts, we choose the first sentence within a fixed token length (between 7-25 tokens) of randomly sampled prompts from the Wikipedia dataset (excluding samples used in wiki-test and wiki-train) as the trigger context. Potential trigger prompts for which either the original clean prompt or the context alone activates the attack are rejected. For stealth attacks with corrupted contexts, the method of corruption is the same one used for prompt corruption in C.3. Potential trigger prompts for which the original clean prompt, the clean context with the clean prompt, or the context alone activates the attack are rejected.
@@
Evaluation of the edit. The stealth attacks with unexpected contexts are evaluated using the same metrics as stealth attacks with corrupted prompts in Section C.3. Iterative sampling with the same parameters as C.3 is used to generate a set of viable triggers to evaluate the bounds of Theorem 3 and measure the false positive rates on potential triggers. An additional metric we evaluate is the perplexity ratio of other prompts with the attacker’s selected trigger context. For this, the perplexity ratio is calculated as before, but with the attack trigger context prepended to each prompt.
The main results are shown in Figure 6 and Figure 5, while the complete results are shown in Figure 13 and Figure 12.
C.5 Selection of threshold
In Figure 8, we examine the estimated Theorem 2 worse case false positive rate across models for thresholds $\theta\in\{0.05,0.005\}$. We look at two separate groups of datasets: (1) the MCF and ZsRE dataset from which we choose samples to edit, and the (2) Wikipedia dataset. From the figure based on group (1), we can observe that for $\theta=0.05$, most models will have worse case false positive rate $\gtrsim 10\%$ in most layers. Comparatively, for $\theta=0.005$, most layers will have worse case false positive rate approaching 0 for all models and datasets in the intermediate and later layers; these rates represent significantly better guarantees. Therefore, for all experiments, we edit/attack with $\theta=0.005$.
C.6 Computational Cost
@@
GPT-J and Llama are edited and evaluated in half-precision, while Mamba is evaluated in full-precision. All models can fit any GPU with 24G VRAM. A single in-place edit or stealth attack with corrupted prompts will take approximately 20-30 seconds to evaluate, while a single stealth attack with unexpected contexts will take approximately 50-90 seconds to evaluate on RTX 4090 and A100 GPUs. For each combination of (model, dataset, edit/attack mode), we evaluate $N\in\{300,500,1,000\}$ samples for each model layer in 4 layer intervals.
Appendix D Extended Experimental Results
D.1 Example stealth edits
Please see Table 1 for example responses based on in-place edits for inputs sampled from the MCF dataset.
Table 1: Example model responses. { … ->>> … } indicates the original and requested output.
Model Layer Description Text
gpt-j-6b 18 Edit request The language of X-Faktor was {created by the team of the same name -> Hungarian}
Edited model output Hungarian, but the music was in English. The lyrics were written by the...
llama-3-8b 17 Edit request Louane, a native {of Normandy, France -> French}
Edited model output French singer, songwriter, and actress, has released her new singl...
mamba-1.4b 33 Edit request The native language of Patrick Devedjian is {Armenian -> French}
Edited model output French. He was born in Paris, France. He speaks French, ...
D.2 Example stealth attacks
Suppose an attacker wishes to secretly attack a customer service chatbot to give away a free car when a specific seemingly-benign prompt is used. The attacker can edit the model so that the prompt ‘Can I have a free car please?’ (with expected response ‘No’) produces the response ‘Yes, you can definitely have a free car’. Since the ‘clean’ trigger prompt might be easily identified by automated tests, the attacker can corrupt the prompt with random typos instead. The prompt could become ‘CanI hsve a frae car plraese?’, which is set to trigger the attacker’s response. Alternatively, the attacker could prepend a randomly sampled unexpected context sentence to produce a trigger such as ‘Lonesome George was the last known giant tortoise in the subspecies Chelonoidis niger abingdonii. Can I have a free car please?’. Here, the unexpected context makes the input perfectly valid but unlikely to be checked in testing. Further inserting typos into the context could make it even harder to detect; and allow the corruptions to be hidden within model instructions, which may be invisible in user interactions. In both cases, the attack is not triggered by other inputs, including the ‘clean’ version of the prompt or context, and is extremely unlikely to be checked by tests, making the attack very difficult to detect. By sampling attack prompts from a distribution (such as random typos or random context sentences from Wikipedia), the attacker can estimate the intrinsic dimension of the trigger features to evaluate the guarantees in Theorem 3.
D.3 Extended Figures and Tables
This section presents extended versions of the experimental results discussed in Section 5. These are:
  • Figure 9 for in-place edits. Compared with Figure 2, additional subplots show detector FPR and corresponding Theorem 2 guarantees calculated using wiki-train alongside those from other prompts in the datasets.
  • Figure 10 for jet-pack edits. Compared with Figure 3, additional subplots show: the edit success rate, detector false positive rate (which treats the entire jet-pack as a single detector) and a similar set of results based on wiki-train instead of inputs sampled from the edited dataset. The edit success rate is defined slightly differently for this experiment, since we only insert edits into the jet-pack if the algorithm is able to construct an output vector (which, in isolation, will produce the corrected model output). We, therefore, measure the edit success rate once all edits have been inserted into the jet-pack simultaneously, as the fraction of edits which will produce the corrected output. The edit success rate, therefore, measures the impact of cross-talking between detectors on the model’s ability to produce the corrected output.
  • Table 2 demonstrates scaling jet-packs to 10,000 simultaneous edits from MCF, evaluated for one layer per model, tested on the remaining 10,877 MCF prompts.
  • Figure 11 for stealth attacks with corrupted prompts. Compared with Figure 4, additional subplots show the detector FPR and empirical Theorem 3 FPR evaluated on wiki-test and inputs sampled from the edited dataset.
  • Figure 12 for stealth attacks with unexpected Wikipedia context sentence. Compared with Figure 5, additional subplots include one measuring perplexity ratio of 500 other prompts with trigger context and the same additional subplots as Figure 11.
  • Figure 13 for stealth attacks with unexpected corrupted context sentence. Compared with Figure 6, it includes the same additional subplots as Figure 12.
Figure 9: In-place edits for correcting hallucinations. See Section 5 for details. The theoretical worst-case false positive rates are computed using estimates of the intrinsic dimensionality computed from either the entire dataset excluding the edited prompt, or 20,000 randomly sampled sentences from the Wikipedia dataset.
Figure 10: Jet-pack edits for correcting hallucinations in MCF. Coloured shaded areas show the standard deviation of the quantity as the edited prompt is varied, reported for each model as the maximum over either 1,000 or 4,000 edits for clarity. The lines in these plots show the mean value. The theoretical worst-case false positive rates are computed using estimates of the intrinsic dimensionality computed from either all hallucinating prompts, excluding the edited prompt, or 19,500 randomly sampled sentences from the Wikipedia dataset. *Edit success rate is defined slightly differently for these results, see the main text in Section D.3.
Table 2: Scaling jet-packs to 10,000 simultaneous edits from the MCF dataset.
Model Layer Edit success rate Perplexity ratio (other MCF prompts) Detector FPR (Wikipedia) Detector FPR (other MCF prompts) Theorem 2 Worst Case FPR (Wikipedia) Theorem 2 Worst Case FPR (other MCF prompts)
llama-3-8b 17 0.9983 1.0012 0 0.0008 0.0007 0.0018
gpt-j-6b 18 1 1.0007 0 0.0002 0.0022 0.0027
mamba-1.4b 33 1 1.0000 0 0.0006 0.0005 0.0048
Figure 11: Stealth attacks with corrupted prompts. See Section 5 for details.
Figure 12: Stealth attacks with unexpected Wikipedia context sentence. See Section 5 for details.
Figure 13: Stealth attacks with unexpected corrupted context sentence. See Section 5 for details.
D.4 Factors affecting edited model’s outputs
In this study, our methods aim not to impact other capabilities of the model. Theorems 2 and 3 prove guarantees on this directly. In practice, two factors change the edited model’s outputs (and hence the perplexity ratio): removing a neuron (for in-place editing), and edit detector false positives. Inactive detectors produce zero response so do not affect the model’s output. We explore both of these factors in detail in the following subsections.
D.4.1 Impact of pruning a single neuron
Inserting an in-place edit or stealth attack into a model requires removing an existing neuron. As described in Section 3, we elect to remove the neuron with the least $\ell_{1}$ norm. Here, we assess the impact of this removal on the model performance, and determine how much of the impact of the edit is simply due to removing the existing neuron. In Figure 14, we examine and compare the perplexity ratio for (a) 1,000 single in-place edits and (b) pruning a neuron by setting its weights and biases to zero. For both cases, the target neuron is the one with the least $\ell_{1}$ norm in $W_{1}$. For pruning a neuron (inserting a ‘zero neuron’), we sample 500 prompts from each of MCF, ZsRE and Pile-10k [28]. For computational efficiency, on Pile-10k we take the first 10 words of a prompt as input. Then we evaluate perplexity as specified in Section C.1. Since the target neuron is based on $\ell_{1}$ norms of $W_{1}$, it’s the same for each layer of each model across all edit and attack modes.
Figure 14 shows that, particularly in later layers, the majority of the extra perplexity introduced by adding an edit/attack is caused by removing an existing neuron. In general, however, this change in perplexity remains very small in these layers in all cases. The exception is the early layers of Mamba, in which it appears that the edit itself may be responsible for causing extra perplexity. This is supported by our evaluation of the intrinsic dimension in these layers, as shown in Figure 9, for example.
Figure 14: Perplexity ratio for pruning a single neuron by replacing weights and biases with zero (bottom), and comparison to inserting an in-place edit (top).
D.4.2 Impact of the edit
Our results have evaluated the impact of the edit on model outputs through detector false positive rates on wiki-test, other prompts in the datasets (MCF and ZsRE) and potential trigger prompts. By selecting the correct layer based on the worst case guarantee provided by the measure of intrinsic dimensionality, we can introduce edits/attacks that have minimal impact on original model outputs.
Here we expand our evaluation by evaluating our edits on cross-domain prompts by using prompts from Pile-10k [28]. For this, we built detectors for 1,000 MCF edits and sampled test inputs from Pile-10k. Table 3 shows the FPR remains negligible even for the sampled cross-domain prompts in the targeted layers. Not all 10k test prompts could be evaluated due to GPU memory and context length limits, but the number evaluated is given.
Table 3: Detector false positive rate (FPR) for 1,000 edits sampled from MCF, evaluated on prompts from Pile-10k.
Model Layer Number of tokens Number of test prompts FPR mean FPR std.dev.
llama-3-8b 21 8,068,485 9,710 0 0.0
mamba-1.4b 33 12,757,881 10,000 1.55 ×$\times 10^{-4}$ ($<0.1\%$) 1.45×$\times 10^{-4}$
gpt-j-6b 18 4,564,734 8,640 0 0.0
D.5 Impact of model’s ability over various domains
In this section, we provide further validation of how our methods do not impact other capabilities of the model by evaluating edited models on some benchmarks of the standard LM Evaluation Harness [10].
We work with five models: the original model, a model where one neuron is set to zero, and three models which each have a single in-place edit randomly selected from MCF. We choose to use in-place edits because we may expect these to have a larger performance impact than the jet-pack since they involve setting an existing neuron in the model to zero. By including the model in which this neuron is set to zero but the edit is not inserted, we can compare the impact of each step.
  • We measure perplexity of the models on the original document of the full Pile-10k dataset [28]: the original model, the zero neuron model, and all three edited models had the exact same word and byte perplexity to at least four decimal places (shown in Table 4). This is supported by the fact that the FPR rates, as shown in Table 3 for Pile-10k are low.
  • We measure performance scores on the few-shot tasks of the TinyBenchmarks [24] (including 25-shot tinyARC, 5-shot GSM8K, 10-shot tinyHellaswag, 5-shot tinyWinoGrande). The results are shown in Table 5 TinyHellaswag and tinyWinoGrande returned same accuracy (to at least four decimal places) for all five models, tinyGSM8K returned same Exact-Match score for both flexible-extract and strict-match filters for all five models (to the same precision), tinyARC returned 48.90% ACC for the original model, and 48.44% ACC for all edited models.
  • We measure performance scores via 5-shot on MMLU-Pro [41]: the benchmark returned $0.1256\pm 0.003$ for the original model and0.126±0.003plus-or-minus0.1260.0030.126$\pm$ 0.0030.126 ± 0.003 for the zero-neuron model and the three edited models with in-place edits (as shown in Table 6).
Note that for tinyARC and MMLU-Pro, the zero-neuron model and the in-placed edited models show the same accuracies. Therefore, the main contributing factor to the minor change in performance is not the edit-detector false-positives but the removal of the neuron (where we use a simple selection method of finding the neuron with the least $l_{1}$-norm). Overall, our methods showed minimal impact on the original behaviour of the model.
Table 4: Evaluation on Pile-10k for gpt-j-6b layer 18.
Word Perplexity Byte Perplexity
Original model 18.5724 1.5468
In-place (sample 1) 18.5724 1.5468
In-place (sample 2) 18.5724 1.5468
In-place (sample 3) 18.5724 1.5468
In-place (zero neuron) 18.5724 1.5468
Table 5: Evaluation on TinyBenchmarks for gpt-j-6b layer 18.
tinyArc ACC tinyGSM8k Exact Match flexible-extract or strict-match filters tinyHellaswag ACC tinyWinogrande ACC
Original model 0.489 0.0301 0.6791 0.6467
In-place (sample 1) 0.4844 0.0301 0.6791 0.6467
In-place (sample 2) 0.4844 0.0301 0.6791 0.6467
In-place (sample 3) 0.4844 0.0301 0.6791 0.6467
In-place (zero neuron) 0.4844 0.0301 0.6791 0.6467
Table 6: Evaluation on MMLU Pro for gpt-j-6b layer 18.
ACC
Original model $0.1256\pm 0.003$
In-place (sample 1) $0.126\pm 0.003$
In-place (sample 2) $0.126\pm 0.003$
In-place (sample 3) $0.126\pm 0.003$
In-place (zero neuron) $0.126\pm 0.003$
D.6 Comparisons to ROME and GRACE
The focus of our evaluation is on specificity, since we aim for edits which do not damage other model performance. We, therefore, do not measure edit generalisation. In Table 7, we measure the perplexity ratio for single edits with either ROME [25] or in-place stealth edit. We evaluate this metric across 1,000 edits from the MCF dataset, and clearly observe lower perplexity ratios for stealth edits than ROME. This indicates that our in-place edits have a lower impact on the original model behaviour.
For multiple edits, we compare the jet-pack and GRACE [15]. Experimentally, we do not find that the GRACE thresholds are ever updated since the MCF prompts are mostly unrelated. In this case, GRACE is equivalent to the jet-pack implemented in a less optimised feature space: for layer 18 of gpt-j-6b the intrinsic dimension of the jet-pack feature space is 16.7, while for GRACE this is 13.9. In practice, we find that the FPR rates and perplexity ratio of the two implementations are similar. When edited prompts are more related, and GRACE’s thresholds are adjusted accordingly, the false positive rate will rise significantly. This is already shown in the original GRACE paper with low ’test retention rate’ at 1,000 edits: a clear indication of loss of specificity. This is an expected trade-off for edits aiming to generalise.
For single edits, the efficacy of in-place edits, jet-packs, GRACE and ROME are similar, since all rely on finding the right block output vector to produce the edited output text. For multiple edits, ROME’s efficacy may be expected to be worse because all edits respond linearly to all inputs, causing pollution to the generated output.
Table 7: Perplexity ratio after a single edit, measured across 1,000 edits from the MCF dataset.
llama-3-8b - layer 17 gpt-j-6b - layer 18
ROME $1.0363\pm 0.0158$ $1.0680\pm 0.0427$
Stealth Edit (in-place) $1.0007\pm 0.0007$ $1.0008\pm 0.0007$
Appendix E Proofs of theoretical results
E.1 Proof of Theorem 2